OpenStudy (anonymous):

what is the formula used in finding the sum of consecutive even integers with an exponent of 2 each?

OpenStudy (anonymous):

S(n) =n(n+1)(2n+1)/6

OpenStudy (agreene):

I don't entirely follow the question but.. n=2k is the even counting function, so I suppose you're asking for: \[\large \sum_{k=1}^{\infty} (2k)^2\]

OpenStudy (anonymous):

@agreene yup. that's what i'm asking for

OpenStudy (anonymous):

S(n) =n(n+1)(2n+1)/6 this is the formula

OpenStudy (anonymous):

S stands for sum :)

OpenStudy (anonymous):

@myko i think that's the formula in finding the sum of the squares of CONSECUTIVE INTEGERS (not even)

OpenStudy (anonymous):

ups, sry didn't read eaven. My bad

OpenStudy (anonymous):

@myko that's alright :D

OpenStudy (kinggeorge):

\[\sum_1^\infty (2k)^2=\sum_1^\infty(4k^2)=4\sum_1^\infty k^2\]From there, you get that the formula is just \[4\cdot {n(n+1)(2n+1) \over 6}\]

OpenStudy (kinggeorge):

The sums I wrote should go to \(n\). Not \(\infty\).

OpenStudy (anonymous):

@KingGeorge thank you so much! :D

OpenStudy (kinggeorge):

You're welcome.